OpenStudy (anonymous):

3x-10>5x+6 answer please

OpenStudy (anonymous):

1) subtract \(3x\) from both sides

OpenStudy (anonymous):

2) subtract \(6\) from both sides

OpenStudy (anonymous):

3) divide both sides by \(2\)

OpenStudy (anonymous):

What's the answer

OpenStudy (anonymous):

lol it is what you get when you do that

OpenStudy (whpalmer4):

It's what you get when you follow those steps correctly!

OpenStudy (whpalmer4):

You're trying to get \(x\) alone one on side of the inequality sign. You can add, subtract, multiply, divide, etc. so long as you do the same thing to both sides of the equation. There's one little wrinkle with inequalities: if you multiply or divide by a negative number, you have to turn the inequality symbol to face the opposite direction. Not an issue in this problem.

OpenStudy (whpalmer4):

Work out an answer and I'll tell you if it is correct or not.

OpenStudy (anonymous):

It's it x=2

OpenStudy (whpalmer4):

Okay, you can't go from an inequality to an equality. Even if you could, that'd still be wrong. Care to share your work? We can find the problem.

OpenStudy (anonymous):

3x-10>5x+6 -3x. -6 3x 4 2x 2. 2 2=x

OpenStudy (whpalmer4):

No. You should write out the entire equation each time, this shortcut of yours isn't helping you get to the correct answer. \[3x-10>5x+6\]subtract \(3x\) from each side: \[3x-3x-10 > 5x-3x+6\]\[-10>2x+6\]Now apply the rest of the steps in the same fashion.

OpenStudy (anonymous):

I keep getting x=2

OpenStudy (whpalmer4):

Show your work! You can't possibly get an equals sign in the answer if you copy the equations as I showed you.

OpenStudy (whpalmer4):

\[-10>2x+6\]Subtract 6 from both sides Show me.

OpenStudy (anonymous):

X>2

OpenStudy (whpalmer4):

Write out the left hand side, and at the end of it, add "-6". Copy the ">" sign. Write out the right hand side. At the end of it, add "-6". Show me the result.

OpenStudy (anonymous):

X>8

OpenStudy (whpalmer4):

What is -10 -6?

OpenStudy (anonymous):

-16

OpenStudy (whpalmer4):

Good, so the left hand side of the equation should be "-16". Now copy the inequality, and you have "-16 >" What is 2x + 6 - 6?

OpenStudy (anonymous):

-8

OpenStudy (whpalmer4):

You don't know the value of x! How are you coming up with a number? Again. What is "2x+6-6"?

OpenStudy (anonymous):

I'm confuse

OpenStudy (whpalmer4):

Simplify the expression "2x+6-6"

OpenStudy (anonymous):

2x

OpenStudy (whpalmer4):

Right. \[-16 > 2x\]Apply the next step, divide both sides by 2

OpenStudy (anonymous):

-8

OpenStudy (whpalmer4):

What is -16/2?

OpenStudy (whpalmer4):

\[\frac{-16}{2}>\frac{2x}2\]Simplify that.

OpenStudy (anonymous):

Is -8

OpenStudy (whpalmer4):

What is -8? Where did the inequality sign go? Where did the \(x\) go?

OpenStudy (anonymous):

X>-8

OpenStudy (whpalmer4):

Keep the stuff in the same order!

OpenStudy (anonymous):

That's the answer

OpenStudy (whpalmer4):

No, it's not! You have to maintain the order of the expressions unless you adjust the inequality sign. What is the simplification of \[\frac{-16}{2}\]

OpenStudy (anonymous):

8/1

OpenStudy (whpalmer4):

No. -16 divided by 2 is not 8. 8*2 does not equal -16.

OpenStudy (anonymous):

-8

OpenStudy (whpalmer4):

That's better. What is the simplification of \[\frac{2x}{2}\]

OpenStudy (anonymous):

1

OpenStudy (whpalmer4):

Where did the \(x\) go?

OpenStudy (anonymous):

X>1

OpenStudy (whpalmer4):

What is the simplification of \[\frac{2x}2\]

OpenStudy (anonymous):

1x

OpenStudy (whpalmer4):

Yes. so what is the simplification of \[\frac{-16}{2} >\frac{2x}{2}\]

OpenStudy (anonymous):

1x

OpenStudy (whpalmer4):

Incredible. What happened to all those numbers and the inequality sign?

OpenStudy (anonymous):

-8>1x

OpenStudy (whpalmer4):

Yes!!! You can also write that as \[-8 > x\]or\[x < -8\] Notice when I put \(x\) on the other side I changed the direction the inequality sign points. \[x < -8\] and \[x > -8\] are NOT equivalent.

OpenStudy (whpalmer4):

The arrow points at the smaller quantity, if you need a memory aid.

OpenStudy (anonymous):

So x >-8

OpenStudy (whpalmer4):

NO!!!!!! Go read my last three posts again.

OpenStudy (anonymous):

X<-8

OpenStudy (whpalmer4):

Yes. Here's the entire problem: \[3x-10>5x+6\]Subtract \(3x\) from each side: \[3x-10-3x > 5x+6-3x\]Collect like terms: \[-10 > 2x+6\]Subtract 6 from each side: \[-10-6 > 2x + 6 - 6\]Collect like terms: \[-16 > 2x\]Divide each side by 8: \[\frac{-16}{2} > \frac{2x}2\]Reduce fractions: \[-8 > x\]Reverse order of inequality to put \(x\) on left side:\[x < -8\] Write that down in your notebook. Study it carefully.